A human cannon ball is shot from a cannon at a 35 degree angle above the horizontal, with an Initial velocity of 36.5 m/s. A net is positioned 115m from the end of the cannon, with the bottom of the net 12m above the opening of the cannon and top of the net 15m above the bottom of the net.


A.) How much time does it take the human cannon ball to to reach the distance the net is at? (Please show work)


B.) Does the human canon ball land in the net or miss it? (Please show work)
Answer #1

V = 36.5 ( cos 35 i + sin 35 j ) = 29.9 i + 20.9355 j


x = 115 m


s = vx t


115 = 29.9 t

t = 3.846 seconds


B) y = 12 m


s = Vy t

Vy t = 3.846 * 20.9355 = 80.517 m


So it lands in the net